7. A surveyor measures the distance across a straight river by the following method (Fig. here given). Starting directly across from a tree on the opposite bank, she walks d = 100 m along the riverbank to establish a baseline. Then she sights across to the tree. The angle from her baseline to the tree is θ = 35.0°. How wide is the river?
1
Expert's answer
2020-11-20T07:17:15-0500


From the definition of the tangent of an angle for a right-angled triangle

"\tg{(C)}=\frac{BA}{BC}"

write down

"BA=BC \cdot\tg{(35^0)}=100 \cdot 0.7=70m"